OpenStudy (polargirl12):

Find the amount of money in an account after 5 years if $8000 is invested at 5 3/4% compounded monthly. Please explain

jimthompson5910 (jim_thompson5910):

You'll use the compound interest formula \[\LARGE A = P*\left(1+\frac{r}{n}\right)^{n*t}\] A = final amount in account after t years P = amount deposited r = interest rate n = compound frequency (how many times you compound interest per year) t = time in years

jimthompson5910 (jim_thompson5910):

In your case A = unknown P = 8000 r = 0.0575 n = 12 t = 5 Things to note: 5 & 3/4% = 5.75% = 5.75/100 = 0.0575 n = 12 because compounding monthly means we compound 12 times a year

OpenStudy (polargirl12):

ok

OpenStudy (polargirl12):

that makes sense

OpenStudy (polargirl12):

so it would be $10657.40?

jimthompson5910 (jim_thompson5910):

`so it would be $10657.40?` correct

OpenStudy (polargirl12):

thank you jim_thompson5910